OpenStudy (anonymous):

Would anyone be so gracious as to helping me get a better understanding of the dreadful algebra!??!?!!? :) Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system. 2x + 9y = 4 3x + 7y = -7

OpenStudy (venomblast):

1st you want to figure out which variable you want to solve

OpenStudy (anonymous):

X

OpenStudy (venomblast):

ok. oh n btw i replied. the answer is A!!! fromt he previous question

OpenStudy (anonymous):

thank you love!!!! :)))))))))))) I got that answer after I "closed" the question box

OpenStudy (venomblast):

i said it 2 times. lol anywayz let do this problem

OpenStudy (anonymous):

thank you nonetheless! ok let's go!

terenzreignz (terenzreignz):

Hey, an intermission, if you would oblige me... if x = y, then ax = ay right?

terenzreignz (terenzreignz):

And if a = b and c = d then a + c = b + d Don't worry, these are relevant, we'll use them.

OpenStudy (anonymous):

Thank you kind sir....! @terenzreignz

OpenStudy (venomblast):

man i hate the ax=ay method

terenzreignz (terenzreignz):

So you have 2x + 9y = 4 From our first rule, we can get 7(2x + 9y) = 7(4) 14x + 63y = 28, let's stop here, and call this "equation 1" You also have 3x + 7y = -7 Again, from our first rule, we can get -9(3x + 7y) = -9(-7) -27x - 63y = 63, stop here again, call this "equation 2" Are you getting it so far?

OpenStudy (anonymous):

Yes, I've gone thru this a time before! But one question, how did you get the "7" in 7(2x+9y)

OpenStudy (anonymous):

If you don't mind me asking.....! :))

OpenStudy (anonymous):

Here's a joke: Subway called, they want their 6 inch back ;)

OpenStudy (venomblast):

LOL

OpenStudy (anonymous):

Another joke/flattery I made up: Verizon called, they want their "SMART" phone back

terenzreignz (terenzreignz):

Ok... You said you wanted to solve for x first (personally, I'd have gone for y first) So, we have to somehow get rid of the y. To do that, you have to make it so that the coefficient (the number that goes beside the letter) of the y in the first equation is equal to the NEGATIVE of the coefficient of the y in the second equation. In the first equation, the coeff of y was "9" and in the second, it was "7" You need to get what's called their "least common multiple" And that's 63. 63/9 is 7, so that's what I multiply to the first equation. 63/7 is 9, so that's what I multiply to the second equation, and then I NEGATED it (multiplied it by -1)

terenzreignz (terenzreignz):

You'll see why this is done, when I proceed... shall I?

OpenStudy (venomblast):

please do lol

OpenStudy (anonymous):

oh my god....this is actually making sense!!!!! :')

terenzreignz (terenzreignz):

Remember our second rule? if a=b and c=d then a+c = b+d ? We'll just apply that, on a larger scale, sort of...

OpenStudy (anonymous):

I remember! I will, I promise!

terenzreignz (terenzreignz):

All right, equation 1: 14x + 63y = 28 equation 2: -27x - 63y = 63 Remember these?

OpenStudy (anonymous):

I do, yes

terenzreignz (terenzreignz):

Now we use the second rule, we add both expressions on the left, and equate them to the sum of the expressions on the right: 14x + 63y - 27x - 63y = 28 + 63

terenzreignz (terenzreignz):

Notice something?

OpenStudy (venomblast):

combine like term *cough**cough*

OpenStudy (anonymous):

I'm writing down the steps to refer to, but yes, I notice that in order to prolong the steps, I'll have to combine

terenzreignz (terenzreignz):

and...?

OpenStudy (anonymous):

Get x by itself

terenzreignz (terenzreignz):

And then you'll have reduced it to a linear equation in one variable. Solve for x, and you're almost done! Just substitute that x which you get for the x in either one of the equations you started with, and then solve for y, and then you'll be done!

OpenStudy (anonymous):

oh my....I appreciate your efforts!!

terenzreignz (terenzreignz):

I appreciate your appreciation; have fun! :)

OpenStudy (anonymous):

I will never have fun with math, but thank you:)